SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Kim Hojeong, Mark Tibbetts.

Operations

  • Hojeong is doing commissioner training - as part of this, she's carrying the pager for the next 2 weeks. Debbie will be keeping an eye on things as back-up.
  • Someone else will have to take the pager 24/8 - 3/9.

EMT

  • See hypernews post about EMT calibration. While the EMC is being kept out of the global calibration, the pilot is requested to specifically include the EMT.
    Should check the calibration actually picks up problems, perhaps during the down.

DCT/DCZ

  • The DCH is seeing more problems again with LV drop-outs in FEA 1. See DCH operations hypernews for details. When they change the FEA we'll do another DCH-DCT calibration.

L3

  • Emailed today because logbook entries were not filled when run 75638 ended with an error. Rainer filled in L3 entries from the hbook files, but the epics variables will have to be filled in from the archiver. It's unclear why the script didn't run.

Trickle injection

  • We need to be able to look at events inside the offline L3 window, in case the trickle injection frequency is changed and we need to change the window size to avoid loosing physics. It can be set up so that Elf makes two passes for every run, one with the window in place and one with the window reversed (so we only process events inside the window).

Trigger Configuration

  • See status page for details of updated analyses. The only inportant study missing is the ISR non-resonant.
  • 1E study is ready, we ought to prepare to take a few runs with the line included in the configuration before the end of the Run. We also ought to be careful not to clash with the SVT and DCH tests of their modified dataflow software (modified to remove deadtime).
  • Arik has been in touch with information pertaining to his study - Hojeong is looking into this.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.